Convert 27 to base 2
Answer by stanbon(57347)

27/2=13 ; remainder 1
13/2=6 ; remainder 1
6/2=3 ; remainder 0
3/2=1 ; remainder 1
1 is the left-most digit of 27(base 2)
27(base 2)= 11011
==========
Cheers,
Stan H.